OpenStudy (anonymous):

Find all solutions to the equation. √3sec θ-2=0 Write your answer in radians in terms of π?

OpenStudy (mertsj):

\[\sec \theta=\frac{2}{\sqrt{3}}\] \[\cos \theta=\frac{\sqrt{3}}{2}\] \[\theta=\frac{\pi}{6}, \frac{11\pi}{6} + 2n \pi\]